.
Glenn and Paula want to dig an irrigation channel for their newest garden. Together it takes them 1.7 hours to dig an appropriately sized irrigation channel. If Paula can dig the channel by herself in 2.5 hours, how long would it take Glenn by himself to dig the channel? If necessary, round to the nearest tenth of an hour.

Answer by stanbon(75887)

Glenn and Paula want to dig an irrigation channel for their newest garden. Together it takes them 1.7 hours to dig an appropriately sized irrigation channel. If Paula can dig the channel by herself in 2.5 hours, how long would it take Glenn by himself to dig the channel? If necessary, round to the nearest tenth of an hour.
----
Together rate: 1/1.7 job/hr
---
Paula rate: 1/2.5 job/hr
----
Glenn rate: 1/x job/hr
-----
Equation:
rate + rate = together rate
1/x + 1/2.5 = 1/1.7
-----
(2.5)(1.7) + 1.7x = 2.5x
------
0.8x = 4.25
---
x = 5.31 hrs (time for Glenn to do the job alone)
=================
Cheers,
Stan H.
================